# From Templates to Natural Language: Generalization Challenges in Instruction-Tuned LLMs for Spatial Reasoning

### v1

@@ Abstract
Instruction-tuned large language models (LLMs) have shown strong performance on a variety of tasks; however, generalizing from synthetic to human-authored instructions in grounded environments remains a challenge for them. In this work, we study generalization challenges in spatial grounding tasks where models interpret and translate instructions for building object arrangements on a $2.5$D grid. We fine-tune LLMs using only synthetic instructions and evaluate their performance on a benchmark dataset containing both synthetic and human-written instructions. Our results reveal that while models generalize well on simple tasks, their performance degrades significantly on more complex tasks. We present a detailed error analysis of the gaps in instruction generalization.
@@ 1 Introduction
In recent years, LLMs have demonstrated impressive capabilities in following natural language instructions across a wide range of tasks (Brown et al., 2020; Yang et al., 2024; Firoozi et al., 2025). However, their ability to generalize from training data to human-authored instructions remains limited in grounded environments, such as collaborative block manipulation (Volovikova et al., 2024), table setting (Shridhar et al., 2020), and structure building (Chalamalasetti et al., 2024; Chaturvedi et al., 2024), where language must be translated into precise spatial actions.
In contrast, studies in human communication (Fischer and Moratz, 2001; Li et al., 2016; Marge et al., 2017) show that people adapt their language depending on the context and the intended partner. Speakers tend to provide step-by-step directions and fewer perspective-dependent expressions when instructing a robot.
Yet despite the simplified instructions, instruction-tuned LLMs often struggle (Kranti et al., 2024) with spatial tasks when evaluated on human-authored language. This raises a question: if human speakers adapt their language to be more robot-friendly, why do models trained on such data still fail to generalize?
@@
For instance, the same speaker might say “put the red block on top of the blue one” in one situation, and “make a staircase” in another. These variations challenge model generalization across the linguistic spectrum.
We investigate this question in the context of structure-building tasks, where instructions guide the manipulation. To better understand the limits of instruction generalization, we fine-tune LLMs (see Figure 1) using synthetic, template-based instructions that reflect the explicit and unambiguous style. We examine whether synthetic-only fine-tuning can enable robust generalization by comparing model performance across tasks that vary in linguistic complexity and spatial structure.
We leverage an existing dataset (Kranti et al., 2024) that provides both synthetic and human-authored instructions for the same grounded spatial reasoning task. Each task involves building specific arrangements of colored shapes on a 2.5D grid (8x8), referred to as “boards.” Every board is annotated with synthetic and human-authored instructions, enabling direct comparison of model behavior across instruction styles in a shared task.
We fine-tune instruction-tuned LLMs using only the synthetic instructions and evaluate their ability to generalize to human-authored ones. Our results show that while models perform well on simple boards, achieving high execution scores, performance drops significantly on regular boards. This suggests that instruction generalization is closely tied to the compositional and structural demands of the underlying task and that fine-tuning only on synthetic datasets is insufficient for robust transfer to more complex spatial concepts.
Our contributions are: (a) we conduct a systematic analysis of generalization to human instructions in structure-building tasks; (b) we show that LLMs fine-tuned on synthetic data perform poorly on human-authored instructions for repetitive tasks.
@@ 2 Related Work
Synthetic vs. Natural Language Generalization: instruction-tuned models often struggle to generalize from synthetic data to natural language (Li et al., 2023; Nwankwo et al., 2025; Shi et al., 2025; Nadas et al., 2025). This highlights the challenges of linguistic complexity in real-world instructions. Our work extends this to the spatial grounding task, showing similar failures in generalization despite high performance on synthetic tasks.
@@
Fine-tuning on Synthetic and Human Instructions in Structure-Building Tasks: prior work has explored fine-tuning LLMs for collaborative structure building in environments like Minecraft (Chalamalasetti et al., 2024; Chaturvedi et al., 2024). These studies typically fine-tune on a mix of synthetic and human-authored instructions but still report limited generalization. Furthermore, Minecraft dataset does not distinguish based on complexity and structures with and without object repetitions, making it difficult to isolate the impact of fine-tuning. Our work builds on these by analyzing performance gaps across instruction styles and board types that affect generalization.
Model In-domain Out-of-domain
Simple Boards Regular Boards Seen Unseen
@@
CodeLLama-7B 0.09 0.32 0.02 0.05 0.56 0.03 0.55 0.04
CodeLlama-34B 0.10 0.26 0.02 0.12 0.16 0.06 0.17 0.07
Qwen2.5-Coder-7B 0.07 0.59 0.00 0.25 0.60 0.00 0.55 0.00
Qwen2.5-Coder-32B 0.14 0.48 0.20 0.23 0.86 0.94 0.82 0.95
Qwen3-8B 0.00 0.54 0.00 0.11 0.00 0.00 0.00 0.00
Qwen3-32B 0.00 0.48 0.00 0.08 0.00 0.00 0.00 0.00
Llama3.3-70B 0.11 0.78 0.07 0.15 0.00 0.16 0.00 0.20
Claude-3 0.43 - 0.25 - - - - -
GPT-4o 0.17 - 0.31 - - - - -
Human baseline 0.98 - 0.76 - - - - 0.95
Table 1: Execution success rates before and after fine-tuning. In-domain: human instructions from the training domain; Out-of-domain: TidyBot dataset: Seen and Unseen sets. Best post-fine-tuning scores are shown in blue.
3 Task Formulation and Datasets
We focus on a spatial reasoning task where the objective is to execute a sequence of spatial actions based on natural language instruction. Our primary dataset is from (Kranti et al., 2024), which involves building a board i.e., arranging objects on a grid. An object is a structured combination of shapes (e.g., washers, screws, nuts, and bridges) arranged in a specific pattern. Each board is categorized by its type: either simple (non-repetitive arrangements) or regular (repetitive arrangements).
Figure 2: An example task illustrating the alignment between the code representation (left), its target structure (top-right), and two instruction styles (bottom).
Each task instance (see Figure 2) consists of an instruction paired with a target code representation, a Python function describing the intended final arrangement on the grid. The goal of the model is to generate this function based on the given instruction and environment context. Furthermore, each task instance is associated with two variants of instructions: (a) synthetic instructions, automatically generated using templated grammar. These are structurally consistent, semantically unambiguous, and direct; (b) human-written instructions collected from annotators (see Appendix A). These use free-form natural language to describe the same goals as the synthetic instructions.
We prompt instruction-tuned language models with both instruction types and evaluate the resulting code by executing it in a virtual simulated environment. The generated board is then compared to the ground truth to compute execution success—the proportion of cases where the model’s output yields the correct spatial configuration.
To test generalization, we fine-tune models on the training set of synthetic instructions and evaluate performance on both synthetic and human-written instructions. This setup allows us to study how well instruction-tuned LLMs generalize from synthetic (templated) instructions to natural human language within the same domain and across tasks. To assess cross-domain generalization, we evaluate these models on the TidyBot (Wu et al., 2023) dataset, which comprises natural language instructions for tidying up real-world household environments. The instances in this dataset are not seen during training (both subsets: seen & unseen).
4 Experimental Setup
We evaluate a diverse set of pre-trained LLMs with varying sizes and architectures, focusing on their ability to interpret grounded spatial instructions and generate executable action sequences. These models are used as-is and fine-tuned on synthetic data only, without exposure to human-written instructions during training.
Models Evaluated
We include both code-centric models (CodeLlama-7B, 34B; Qwen2.5-Coder-7B, 32B) and general-purpose models (Qwen3-8B, 32B; LLaMA-3.3-70B). All models are loaded in 4-bit precision using the Unsloth library [^1] https://unsloth.ai/. We compared these results against the reported scores of the closed models GPT-4o and Claude-3 (version-opus) by Kranti et al. (2024).
Fine-Tuning
Models are fine-tuned on synthetic instructions paired with target code representations. The training set comprises a mixture of Simple Boards (1072 boards) and Regular Boards (1168 boards) (Kranti et al., 2024). We adopt a chat-style prompt format, including environment details (more details in Appendix A.1), followed by the instruction and target code. We use the QLoRA configuration provided by Unsloth with the following hyperparameters: 1 epoch, 2 steps, adam-8bit optimizer, rank $r=16$, lora_alpha=16, lora_dropout=0.0, batch size=2, learning rate=$2\mathrm{e}{-4}$.
Evaluation Metrics
We assess model performance on instruction-following capability, execution accuracy, and cross-domain robustness. Our goal is to evaluate whether models can generalize across instruction types (synthetic $\rightarrow$ human) and across domains (arranging board $\rightarrow$ sorting objects).
Error Rate: A metric capturing instruction adherence based on predefined response constraints (output formatting) in the prompt (Appendix A.1).
Execution Success Rate: Our primary task success metric. A prediction is successful if the board configuration in the generated response matches the gold sequence when executed in a deterministic simulator. This involves matching correct component usage, spatial placement, and ordering.
Figure 3: Error analysis on Simple Boards (top) and Regular Boards (bottom), with outputs from specific models indicated above each result. Errors highlight challenges in spatial generalization from natural language instructions.
@@ 5 Results
The results are presented in Table 1. We present results for (a) generalization from synthetic to human instructions on the structure-building dataset and (b) transfer to human-written instructions in the TidyBot dataset (Wu et al., 2023).
5.1 In-Domain Generalization: Synthetic → Human
Table 1 reports execution success, and Table 5 (in Appendix A.2.3) reports error rates on simple and regular boards for each model for human instructions and the findings are: (a) fine-tuning substantially improves success on human instructions, particularly for models like Qwen2.5-7B, Qwen2.5-32B, and LLaMA-3.3-70B; (b) models perform better on simple boards than on regular boards, indicating that object repetition (and resulting referential ambiguity) challenges generalization; (c) large models such as GPT-4 and Claude-3, evaluated few-shot, show moderate ability to generalize to human instructions but underperform compared to fine-tuned models. High human baseline scores indicate that the task has a clear upper bound, which current models have yet to reach.
5.2 Cross-Domain Generalization: Transfer to TidyBot
To test whether models generalize beyond vocabularies seen during training, we evaluate models fine-tuned on synthetic structured boards on another domain dataset: the TidyBot dataset (Wu et al., 2023), which contains free-form human instructions for real-world object arrangement. The results are reported in Table 1 and the findings are: (a) top-performing models (Qwen2.5-32B, LLaMA-3.3-70B) maintain high success show generalization to unseen instructions and unseen objects, suggesting that fine-tuning on well-structured, synthetic spatial tasks can yield transferable skills; (b) other models (e.g., Qwen2.5-7B, CodeLlama-7B) degrade significantly, indicating possible overfitting.
Fine-tuning on synthetic instructions enables strong generalization to human-written instructions in low-ambiguity environments. Referential ambiguity (see Appendix A.3), introduced by repetitive objects, significantly correlates with failures. Some models learn instruction-to-structure mappings that generalize across language form (synthetic → human) and object domain (seen → unseen).
5.3 Qualitative Error Analysis
We categorize incorrect executions across models by error type (see Figure 3). The analysis indicates: (a) Element mismatch is most common on synthetic instructions; (b) Hierarchical instructions and ambiguous references dominate human instructions in regular boards, reinforcing the idea that repetition introduces referential entropy; (c) Models like Llama-3.3-70B and Qwen2.5-32B exhibit fewer invalid sequences, suggesting that fine-tuning improves overall generalization.

### v2

@@ Abstract
Instruction-tuned large language models (LLMs) have shown strong performance on a variety of tasks; however, generalizing from synthetic to human-authored instructions in grounded environments remains a challenge for them. In this work, we study generalization challenges in spatial grounding tasks where models interpret and translate instructions for building object arrangements on a 2.52.5D grid. We fine-tune LLMs using only synthetic instructions and evaluate their performance on a benchmark dataset containing both synthetic and human-authored instructions. Our results reveal that while models generalize well on simple tasks, their performance degrades significantly on more complex tasks. We present a detailed error analysis of the gaps in instruction generalization.
@@ 1 Introduction
Accurate spatial grounding (Green et al., 2006; Brenner, 2007) is important for effective human-robot interaction (Bisk et al., 2016; Shridhar and Hsu, 2018; Hatori et al., 2018). It involves the robot interpreting spatial references (Hüttenrauch et al., 2006) and relational cues expressed in natural language instructions (Dang-Vu et al., 2015; Paul et al., 2018; Tellex et al., 2020; Bisk et al., 2020). Large Language Models (LLMs) are increasingly used in robotics pipelines (Ichter et al., 2022; Wang et al., 2024b; Lim et al., 2024; Macaluso et al., 2024) to translate such instructions into high-level plans and then into low-level robot actions. While these models have shown progress in instruction-following tasks, they continue to face challenges (Liu et al., 2023; Ding et al., 2023; Jiang et al., 2025) when grounding spatial language references.
Human instructions are often abstract (Minsky, 1980), underspecified (Hendriks et al., 2014), and context-dependent. For example, a user might say “put the red block on top of that blue one” in one case, and “make a staircase” in another. These variations rely on shared context and conceptual inference, making them harder to interpret (Kranti et al., 2024a; Chaturvedi et al., 2024) than synthetic or templated instructions, which are designed to be direct and explicit.
@@
To explore these issues in a controlled setting, we investigate them in the context of a structure-building task, where instructions guide the manipulation of objects on a grid. We leverage an existing dataset (Kranti et al., 2024b) that includes both synthetic and human-authored instructions for the same grounded spatial reasoning task. Each task involves building specific arrangements of colored shapes on an 2.5D grid with size $8\times 8$, referred to as “boards.” Each board is paired with both synthetic and human-authored instructions, enabling direct comparison of model behavior (see Figure 1) across instruction styles within a shared task space.
We evaluate how well pre-trained LLMs perform with both synthetic and human-authored instructions across different board types. Furthermore, we fine-tune instruction-tuned LLMs using only the synthetic instructions and evaluate them whether they generalize to human-authored ones. Our results show that while fine-tuning improves model performance, and models achieve high execution scores on simple boards, performance gains are limited on regular boards. This suggests that instruction generalization is closely tied to the compositional and structural complexity of the task, and that fine-tuning solely on synthetic data is insufficient for robust transfer to instructions that require more abstract or relational understanding.
Our contributions are as follows: (a) we evaluate pre-trained and fine-tuned LLMs on both synthetic and human-authored instructions, and report quantitative performance across different board types (simple vs. regular); (b) we conduct a detailed qualitative analysis of model responses before and after fine-tuning, categorizing the types of errors across instruction styles and board complexity; (c) we analyze the embedding similarity of human-authored instructions, to their synthetic counterparts, the number of shapes involved, and how these factors correlate with model performance; and (d) we test the generalization of fine-tuned models on two related, out-of-domain code generation tasks, and report the performance related to the generalization.
@@ 2 Related Work
Spatial Grounding with LLMs
Spatial grounding refers to the task of interpreting spatial references in natural language and linking them to a target environment. Recent work has explored the use of LLMs for tasks such as block manipulation (Valmeekam et al., 2023; Goldberg et al., 2024), table setting (Liang et al., 2023; Vemprala et al., 2024), tool usage (Xu et al., 2023) and collaborative assembly (Lim et al., 2024; Macaluso et al., 2024; Joglekar et al., 2024). Our work is inspired by these efforts but differs in how the grounding is performed. Prior approaches generate sequences of atomic actions or low-level code that are directly executable in the environment. In contrast, we focus on generating abstract functions that encode spatial semantics but are not immediately grounded; these functions are interpreted and executed by a downstream system. This abstraction allows us to analyze how well LLMs encode spatial structure independently of immediate action execution.
Instruction-to-Code Translation for Spatial Tasks
Instruction-to-code translation focuses on converting natural language inputs into executable programs or structured representations (Singh et al., 2023; Huang et al., 2023a, b; Wang et al., 2024a; Hu et al., 2024). Our work extends these efforts by comparing model performance across two instruction styles, synthetic and human-authored, for the same spatial task. This setup enables us to study generalization and instruction-following behavior under consistent task conditions.
Generalization from Synthetic to Human Instruction
@@
Instruction-tuned models often struggle to generalize well from synthetic to natural language (Li et al., 2023; Nwankwo et al., 2025; Shi et al., 2025; Nadas et al., 2025). This reflects the challenges of handling variation in human-authored instructions. We extend this to a spatial grounding task and report similar failures in generalization despite high performance on synthetic tasks.
Fine-tuning on Synthetic vs. Human Instructions
@@
Prior work has explored fine-tuning LLMs for collaborative structure building in environments like Minecraft (Kranti et al., 2024a; Chaturvedi et al., 2024). These studies typically fine-tune on a mix of synthetic and human-authored instructions and report limited generalization. Furthermore, Minecraft dataset (Narayan-Chen et al., 2019) does not distinguish based on complexity and structures with and without object repetitions, making it difficult to isolate the impact of fine-tuning. Our work builds on these by analyzing performance gaps across instruction styles and board types that affect generalization.
Figure 2: A simple board example showing how code (left) maps to a target structure (right), with both synthetic and human-authored instructions describing the same configuration. Models are expected to generate both the function definition and its usage based on the given instruction.
3 Task Formulation and Datasets
Figure 3: A regular board example illustrating repeated structural patterns defined by code (left), their corresponding target structure (right), and matching synthetic and human-authored instructions. The function definition is provided in the prompt; models are expected to generate only the usage code (e.g., nested for loops) based on the instruction.
Task: it is framed as a two-player game between a programmer and a cobot (collaborative robot), where there is no intermediate feedback and execution happens only at the end of the interaction. The programmer (instruction giver) is given a target board and instructs the cobot (instruction follower) to translate the instruction into code. The code generated by the cobot is executed in a virtual simulation environment and evaluated for correctness.
Boards: We use the SARTCo (Kranti et al., 2024b) dataset for finetuning and the main testbed for our experiments. It consists of $260$ human-authored instructions for a structure building task on a 2.52.5111A 22D grid with stacking support in each cell and without the complexity of full 33D simulation. grid. Each instance in the dataset includes a triplet: a code representation, a target board, and a natural language instruction. An example consisting of the target board, an abstract Python function, and two styles of instructions, is shown in Figure 2 and Figure 3. Each target board is a specific arrangement of objects on a grid. An object is a composition of shapes such as washers, screws, nuts, and bridges. Boards are labeled as either simple, with non-repetitive arrangements, or regular, with repetitive arrangements. The goal is to generate this semantically grounded function from the instruction. This setup reflects instruction-following scenarios and enables evaluation of how well LLMs can spatially ground such instructions.
Instructions: they are of two types: synthetic & Human-authored. Synthetic instructions are generated using a template-based grammar. They are structurally consistent and semantically explicit. They are available in both single-turn and multi-turn formats. Human-authored instructions are free-form language inputs for the same target structure. They are available only in single-turn format. Both types are aligned to the same structure building task.
Learning on Synthetic Instructions: we prompt LLMs with both instruction types and evaluate the performance by executing the generated code in a virtual simulation. The output is compared to the target to compute execution success, defined as the proportion of cases where the model output matches the target board. We fine-tune models on synthetic instructions and evaluate them on both synthetic and human-authored instructions. This setup tests how well models generalize from synthetic to human language in the same task domain.
Cross-domain Datasets: We also evaluate cross-domain generalization using the HEXAGONS (Lachmy et al., 2022) and TidyBot (Wu et al., 2023) datasets. These datasets help assessing whether fine-tuned models are able to perform on different but related spatial instruction tasks.
4 Experimental Setup
We evaluate a diverse set of LLMs with varying sizes and architectures, focusing on their ability to interpret grounded spatial instructions and generate executable action sequences. These models are used as-is (few-shot) and fine-tuned on synthetic data only, without exposure to human-authored instructions during training. We use the clembench (Chalamalasetti et al., 2023) framework to manage player interactions.
4.1 Evaluated Models & Parameters
Models: We include both code-centric models (Qwen2.5-Coder-7B, 32B) and general-purpose, instruction-following models (Qwen3-8B, 32B; LLaMA-3.1-8B, 3.3-70B). All models are loaded in 4-bit quantized precision using the Unsloth library [^2] https://unsloth.ai/. We compared these models against the commercial ones: GPT-4o and Claude-4-Sonnet.
[…]
Hyperparameter Selection and Ablations: To determine optimal fine-tuning settings, we conducted a comprehensive ablation study varying data composition, ordering, and hyperparameters. We evaluated models trained on (i) only simple boards, (ii) only regular boards, and (iii) combined boards with/without shuffling. Additionally, we explored the impact of training sample sizes ($100$-to-$1000$ per board type), batch sizes (22-to-88), learning rates ($1\mathrm{e}{-4}$, $2\mathrm{e}{-4}$), epoch counts (11-to-55), lora_dropout ($0.10$, $0.15$, $0.20$) and early stopping patience($2$, 33). These studies informed our final configuration: combined dataset with shuffled samples, full training set, batch size 8, 3 epochs, learning rate $1\mathrm{e}{-4}$, and dropout $0.10$. More details and ablation study results are available in Appendix A.3.1.
4.2 Evaluation Metrics
We assess model performance on instruction-following capability, execution accuracy, and cross-domain robustness.
Error Rate:
A metric capturing instruction adherence based on predefined response constraints (output formatting) in the prompt (Appendix A.1). In the clembench setup, responses that violate these constraints are marked as abort.
Execution Success Rate:
Our primary task success metric. Each model-generated function is executed in a virtual simulator and compared against the target board (an $8\times 8$ grid with shape type, color, position, and stacking order). A prediction is successful if the resulting board configuration exactly matches the ground truth in component type, color, spatial placement, and stacking sequence.
Human Baseline:
To contextualize model performance, we include a human-generated baseline. We developed a simple user interface (see Figure 10 in Appendix) that shows the human-authored instructions (by a different annotator) on the left and the setup to reconstruct the target board on the right. We hired an annotator and asked to reconstruct target boards. These outputs were executed using the same virtual simulator and computed the same Execution Success Rate metric, allowing direct comparison between model and human performance.
Cross-domain robustness:
We measure generalization by comparing model performance before and after fine-tuning. Specifically, we evaluate whether models trained on synthetic instructions generalize to (1) human-authored instructions in the same task domain (structure building: synthetic $\rightarrow$ human), and (2) structurally similar but domain-shifted tasks such as drawing hexagons (HEXAGONS) and sorting objects (TidyBot).
Model Abort Rate $\downarrow$ Performance $\uparrow$
SB RB SB RB
@@ 4.2 Evaluation Metrics
Qwen2.5-Coder-7B $0.00$ $0.00$ $1.00$ $0.00$ $0.07$ $0.40$ $0.07$ $0.35$
Llama3.1-8B $0.00$ $0.11$ $0.00$ $0.09$ $0.00$ $0.08$ $0.00$ $0.21$
Qwen3-8B $1.00$ $0.38$ $0.95$ $0.59$ $0.00$ $0.09$ $0.00$ $0.12$
Qwen2.5-Coder-32B $0.00$ $0.00$ $0.00$ $0.00$ $0.14$ $\mathbf{0.68}$ $0.23$ $\mathbf{0.54}$
Qwen3-32B $0.95$ $0.00$ $0.97$ $0.00$ $0.00$ $0.53$ $0.00$ $0.38$
Llama3.3-70B $0.00$ $0.00$ $0.72$ $0.06$ $0.11$ $\mathbf{0.69}$ $0.07$ $0.33$
GPT-4o $0.00$ - $0.00$ - $0.72$ - $0.55$ -
Claude-4(sonnet) $0.00$ - $0.00$ - $0.88$ - $0.60$ -
Human-Baseline - - - - 0.98 0.76
Table 1: Abort rate and task performance (accuracy) on simple boards (SB) and regular boards (RB) from human-authored set, before and after fine-tuning on the synthetic set. While some models show gains on SB, improvements on RB are low.
@@ 5 Results
We present quantitative and qualitative results, followed by cross-domain generalization analysis. Quantitative results are organized first by board type (simple and regular) for human-authored instructions, and then by instruction type (synthetic and human-authored) for overall comparison.
Model FSG FSC
SB RB SB RB
Qwen2.5-Coder-7B $0.25$ $0.25$ $0.23$ $0.19$
Llama3.1-8B $0.02$ $0.28$ $0.11$ $0.29$
Qwen3-8B $0.05$ $0.10$ $0.05$ $0.29$
Qwen2.5-Coder-32B $0.45$ $\mathbf{0.53}$ $0.32$ $\mathbf{0.51}$
Qwen3-32B $0.35$ $0.24$ $0.42$ $0.54$
Llama3.3-70B $\mathbf{0.54}$ $0.32$ $\mathbf{0.55}$ $0.15$
Table 2: Model performance with RB prompt styles (FSG: Function Signature, FSC: Function Schematic), evaluated on SB (simple) and RB (regular) boards.
5.1 Quantitative Analysis
Performance based on Board Type
Table 1 presents model accuracy on simple and regular boards for human-authored instructions. Most models perform better on simple boards. Llama3.3-70B and Qwen2.5-Coder-32B achieve higher scores of $0.69$ and $0.68$, respectively. While the scores also improve for regular boards, the gains are smaller. Qwen2.5-Coder-32B achieves $0.54$ accuracy. LLaMA3.1-8B and Qwen3-8B have lower scores, primarily due to hallucinations or failure to follow the expected response format, resulting in aborted executions (see abort rate in Table 1). Although fine-tuning improves accuracy, the gap with the human upper bound remains substantial.
Smaller models (Llama3.1-8B, Qwen3-8B) perform better on regular boards than on simple boards. This may be due to differences in the expected output structure. For regular boards, the target code typically consists of nested for loops that repeat the same object placement across multiple locations (see Figure 3). In contrast, simple boards often require placing a sequence of different shapes with precise spatial relationships, typically implemented using abstract function definitions. This output structure appears to be harder for these models to generate reliably.
Model FSG FSC
SB RB SB RB
Qwen2.5-Coder-7B $0.00$ $0.00$ $0.23$ $0.19$
Llama3.1-8B $0.01$ $0.00$ $0.31$ $0.00$
Qwen3-8B $0.65$ $0.51$ $0.55$ $0.02$
Qwen2.5-Coder-32B $0.00$ $0.00$ $0.00$ $0.00$
Qwen3-32B $0.00$ $0.00$ $0.01$ $0.00$
Llama3.3-70B $0.00$ $0.00$ $0.00$ $0.52$
Table 3: Effect of RB prompt styles (FSG: Function Signature, FSC: Function Schematic) on abort rates for SB (simple) and RB (regular) boards.
Effect of prompt variation
Additional experiments were conducted using different prompt styles for regular boards. Table 2 shows results from two styles: Function Signature (FSG), which includes only the function signature in the prompt, and Function Schematic (FSC), which includes the signature along with a brief schematic description of the function (see Appendix A.2 for more details). These prompt variations do not significantly affect performance on regular boards, but they reduce performance on simple boards. For example, Qwen2.5-Coder-32B drops from $0.68$ to $0.45$ (FSG) and $0.32$ (FSC), while LLaMA3-70B drops from $0.69$ to $0.54$ (FSG) and $0.55$ (FSC). These results suggest that while FSG and FSC are sufficient to support symbolic reuse (as in regular boards), they are less effective when the model is required to generate symbolic abstractions, as in simple boards.
Model SB RB
ST HA ST HA
Qwen2.5-Coder-7B 0.98 0.40 1.00 0.35
Llama3.1-8B 0.46 0.08 0.75 0.21
Qwen3-8B 0.71 0.09 0.65 0.12
Qwen2.5-Coder-32B 0.98 0.68 1.00 0.54
Qwen3-32B 0.99 0.53 1.00 0.38
Llama3.3-70B 0.98 0.68 1.00 0.33
GPT-4o 0.78 0.72 0.64 0.55
Claude-4 0.95 0.88 0.98 0.60
Human-Baseline 0.98 0.76
Table 4: Performance by instruction type: ST (synthetic) and HA (human-authored) on SB (simple) and RB (regular) boards.
Figure 4: Execution analysis of different models on an instruction-to-code generation task. Given a natural language instruction (top left), only Qwen2.5-32B and Qwen3-32B generate correct and executable code that replicates the target board (top right). Other models fail due to issues such as undefined variables, incorrect coordinates, or invalid placements.
Few-shot prompting
We also experimented with few-shot prompting using the fine-tuned models. Table 7 (in Appendix A.2.1) shows that few-shot prompting improves Qwen2.5-Coder-32B’s score on simple boards to $0.82$, but causes a slight decrease in RB performance, from $0.54$ to $0.51$.
Performance based on Instruction Type
We compare model performance on synthetic and human-authored instructions to quantify the generalization gap. Table 4 shows that while models achieve high accuracy on synthetic instructions, performance drops significantly on human-authored inputs. This discrepancy highlights the challenge of transferring symbolic reasoning skills to more abstract and natural instruction styles.
The improved regular board performance seen in Llama3.1-8B, Qwen3-8B may stem from the relative simplicity of generating repetitive structure, particularly when trained on synthetic instructions that explicitly describe iteration patterns.
Overall, fine-tuning on synthetic data improves performance on regular boards across all models. For simple boards, the impact is varied: some models, particularly instruction-tuned or code-oriented ones, benefit substantially, while others show little improvement due to format violations and hallucinations. These findings indicate that even a small amount of synthetic supervision can be beneficial, but the downstream effects depend on model alignment, size, and prompt structure.
5.2 Qualitative Analysis
Figure 4 presents a qualitative comparison of model predictions for a human-authored instruction. The instruction requires placing a green washer in the bottom left corner of a grid, adding a yellow washer next to it on the right, and stacking a red horizontal bridge on top of the two washers. The corresponding ground truth code places the shapes starting from (7, 0) using relative spatial offsets.
Qwen2.5-32B and Qwen3-32B generated code that exactly matched the target representation. These models correctly interpreted the spatial reference: bottom left as (7, 0), maintained the ordering of shape placements, and generated a semantically translated abstract function.
Llama3-70B generated syntactically well-formed code with appropriate functional abstraction, but misinterpreted the spatial reference, mapping “bottom left” to (1, 0) rather than (7, 0). Llama3-8B shows similar behavior of correct abstraction but incorrectly uses symbolic variables (as x, y) for cell positions. Qwen2.5-7B generated correct structure but applied incorrect relative positions, leading to an execution-time placement error. Qwen3-8B generated an unrelated response, indicating failure in instruction following. This suggests that while some models are able to replicate correctly, others struggle with grounding language in spatial coordinates.
Error Categorization
Each model-generated code was executed, and errors were grouped based on either execution failed or succeeded with a semantic mismatch. Responses that failed to execute were classified as board placement errors and for successfully executed code, we compare the resulting board with the ground truth and identify mismatches in shape type, color, or placement order. These are classified as element mismatches.
Figure 5: Error distribution on SB (simple) boards for the Qwen2.5-Coder-32B model after fine-tuning. The primary pie chart shows the proportion of overall error types, with “Board Placement” accounting for 41.5% and 1 “Element. Mismatch” 58.5%. A secondary pie further decomposes the Board Placement errors into Dimensions mismatch (51.9%), Depth Mismatch (25.9%), and Value Error (22.2%)..
Within board placement errors, we observed different error categories (see Figure 5 and Figure 6). DepthMismatch occurs when a bridge is placed without a supporting shape beneath it, typically due to earlier placement errors. BridgePlacement errors involve stacking bridges above the allowed two-level height limit. DimensionMismatch and ValueError arise when object placements exceed board boundaries or use invalid coordinates. Other errors, such as NameError and KeyError, result from the use of undefined variables or incorrect dictionary access (see Figure 4). Additionally, environment-specific errors such as NotOnTopOfScrew and SameShapeStacking reflect violations of symbolic constraints defined by the board logic.
Overall, fine-tuning reduced board placement errors across models, with most remaining errors attributable to element mismatches. (more details are in Appendix A.5).
Instruction Similarity
Table 1 reports improved model performance on human-authored instructions following fine-tuning, with larger gains observed on simple boards and smaller gains on regular boards. This difference may arise from the phrasing of regular board instructions, which often omit explicit spatial details, making it harder for models to generate the intended code. In contrast, instructions for simple boards are more concrete, even when the corresponding code is complex.
To examine the effect of instruction variation, we compute embedding similarity between synthetic and human-authored instructions for the same board configurations. We use BLEU to measure surface-level overlap and Sentence Transformers (Reimers and Gurevych, 2019) cosine similarity to capture semantic alignment.
Figure 6: Error distribution on RB (regular) boards for the Qwen2.5-Coder-32B model after fine-tuning.
Table 5 presents median and range values across board types. Instructions for simple boards show higher embedding similarity between synthetic and human-authored variants and are associated with larger execution accuracy gains. For regular boards, the similarity is lower and performance drops. To further analyze this effect, we examine how similarity varies with the number of shapes in the object. As shown in Table 9 (in Appendix), instructions referring to objects with fewer shapes tend to have greater execution success. These results suggest that lower semantic and lexical alignment between synthetic and human-authored instructions, particularly in structurally repetitive but abstract cases, limits the generalization ability of models fine-tuned only on synthetic data.
Score Ranges BLEU ES
SB RB SB RB
Median Values 0.356 0.024 0.979 0.623
Minimum Values 0.21 0.01 0.92 0.44
Maximum Values 0.66 0.1 0.99 0.82
5.3 Cross-Domain Generalization
To assess whether models generalize beyond the vocabulary and structure observed during training, we evaluate models fine-tuned on synthetic data on other datasets.
Transfer to Hexagons
The HEXAGONS dataset (Lachmy et al., 2022) consists of human-authored instructions for coloring specific cells on a hexagonal grid, based on natural language descriptions. These instructions are similar to repetitive patterns, resembling the regular board structures in our setup. The coloring operation, denoted as paint(color, row, column), is analogous to our put(shape, color, row, column) function. The dataset uses a $10\times 18$ hexagonal grid. For evaluation, we use the test split [^3] https://github.com/OnlpLab/Hexagons/blob/main/data/test.jsonl, which includes $62$ drawing procedures. Table 6 reports model performance on this task before and after fine-tuning. The fine-tuned models show no improvement compared to pretrained models. This limited transferability may be attributed to the domain shift between the training data and the HEXAGONS instructions, which differ in both linguistic style and task structure. Additionally, the larger grid size ($10\times 18$) significantly increases the complexity of spatial reasoning and the difficulty of learning accurate coordinate mappings.
Figure 7: Example from the drawing hexagons task with a human-authored instruction and its corresponding target code.
Transfer to TidyBot
The TidyBot dataset (Wu et al., 2023) contains free-form human instructions for real-world object arrangement. The core operation, expressed as pick_and_place(item, newposition), is similar to shape placement in our setting through put(shape, color, row, column). Results are presented in Table 6. Top-performing models such as Qwen2.5-32B and LLaMA3-70B maintain high task success, demonstrating generalization to other domain instructions and object categories. This suggests that fine-tuning on structured synthetic spatial tasks supports transfer to some domains. In contrast, models such as Llama3.1-8B, Qwen3-8B show significant performance degradation, indicating limited generalization and potential overfitting to the training domain.
Figure 8: Example task from Tidybot with human-authored instruction and corresponding target code.
Model HEXAGONS TIDYBOT
Before After Before After
Qwen2.5-Coder-7B $0.00$ $0.03$ $0.00$ $0.00$
Llama3.1-8B $0.00$ $0.00$ $0.00$ $0.039$
Qwen3-8B $0.00$ $0.00$ $0.00$ $0.00$
Qwen2.5-Coder-32B $\mathbf{0.10}$ $\mathbf{0.11}$ $\mathbf{0.82}$ $\mathbf{0.90}$
Qwen3-32B $0.00$ $0.03$ $0.00$ $0.00$
Llama3.3-70B $0.06$ $\mathbf{0.10}$ $0.00$ $0.06$
Human-Baseline 0.76 0.95
Table 6: Model performance on HEXAGONS and TIDYBOT tasks before and after fine-tuning.